PHY2010-002 Intro to Classical 《Homework 4 CH-6 Constants 1 Periodic Table Youre driving your new sports car at 70 mph over the top of a hill that has a radius of curvature of 540 m mg Provide Feedback 名。 Please solve and explain
Answer #1

speed of vehicle = 70 mph = 31.29 m/s
radius = 540 m

centrifugal acceleration of car= v2/R=31.292/540=1.81 m/s2

when the car is at crest,the weight of car and centrifugal force will act in opposite directions

acceleration due to gravity = 9.8 m/s2

apparanet weight of car = m*(9.8-1.81)=m*7.99 (where m is mass of car)

actual weight of car = m*9.8

required fraction = 7.99m/9.8m=0.815 = 81.5%

This is the required answer